OpenStudy (anonymous):

A polyhedron has 7 vertices and 13 edges. How many faces does it have? Answer A. 6 B. 7 C. 8 D. 9

OpenStudy (eyust707):

\[F + V - E = 2\]

OpenStudy (anonymous):

Use Euler's formula F - E + V = 2, where F is the number of faces, E is the number of edges, and V is the number of vertices of the polyhedron: F - E + V = 2

OpenStudy (eyust707):

V = 7 E = 13

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

F - 13 + 7 = 2

OpenStudy (anonymous):

8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks